Assume that Zimbabwe and Portugal can switch between producing toothbrushes and producing hairbrushes at a constant rate. Machine Minutes Needed to Make 1 Toothbrush Hairbrush Zimbabwe 3 10 Portugal 5 efer to Table 3-6. Assume that Zimbabwe and Portugal each has 180 machine minutes available. If each country divides its time equally tween the production of toothbrushes and hairbrushes, then total production is D a. 48 toothbrushes and 24 hairbrushes. D b. 720 toothbrushes and 1440 hairbrushes. D c. 96 toothbrushes and 48 hairbrushes. O d. 24 toothbrushes and 12 hairbrushes.
Answer #1

Given zimbabwe and portugal each country has=180min

Each country divided time equally between production of toothbrushes and hairburshes i.e=90min

Production ofToothbrush of zimbabwe and portugal=90/3+90/5

=30+18=48

Production of Hairburshes of zimbabwe and portugal=90/10+90/6

=9+15=24

Total production =48tooth production and 24 hairbrushes

Therefore option is a.